OpenStudy (anonymous):

The lengths of the sides of a triangle are represented by three consecutive integers. If the perimeter of the triangle is 96 feet, find the lengths of its sides.

OpenStudy (anonymous):

x+(x+1) + (x+2) = 96. Can you solve now?

OpenStudy (jiteshmeghwal9):

Let the three consecutive sides be (x),(x+1),(x+2).then x+x+1+x+2=96 3x+3=96 3x=96-3 3x=93 x=93/3=31

OpenStudy (anonymous):

thank youu! that really helped.! :D

OpenStudy (anonymous):

You are welcome..:)

OpenStudy (jiteshmeghwal9):

yw:)